# IGNN-Solver: A Graph Neural Solver for Implicit Graph Neural Networks

### v1

@@ 3 Preliminaries
For training parameters $\bm{\theta}=\{{\bm{W}},\bm{\Omega}\}$ during the back-propagation process of neural networks, IGNNs compute the Jacobian $\nabla_{\bm{Z}}\mathcal{L}$ by solving the following equilibrium equation
$\nabla_{\bm{Z}}\mathcal{L}={\bm{D}}\odot({\bm{W}}^{\top}\nabla_{\bm{Z}}\mathcal{L}\hat{{\bm{A}}}^{\top}+\nabla_{\bm{X}}\mathcal{L}),$ (3)
@@
where ${\bm{D}}=\phi^{\prime}({\bm{W}}{\bm{X}}\hat{{\bm{A}}}+b_{\Omega}({\bm{U}}))$ and $\phi^{\prime}(\cdot)$ refers to the element-wise derivative of the map $\phi$. Once $\nabla_{\bm{Z}}\mathcal{L}$ is obtained, we can use the chain rule and the implicit function theorem [28] to easily compute $\nabla_{\bm{W}}\mathcal{L}$ and $\nabla_{\bm{\Omega}}\mathcal{L}$.
@@
In summary, IGNNs usually require numerous iterations to obtain equilibrium, resulting in significant overhead during training and inference. Therefore, it is crucial to find a faster and more efficient method for solving the fixed-point equation of IGNNs.
Figure 1: The overall architecture of the IGNN-Solver compared with the generic Anderson solver at each iteration.
4 Implicit graph neural networks Solver
@@
Although traditional fixed-point solvers for IGNNs (as discussed in Section 1 and  2) demonstrate functionality, they are characterized by relatively slow performance, necessitate manual parameter tuning, and are not well-suited for graph-based applications. To address these limitations, we have designed a lightweight, learnable, and content-aware fixed-point solver for IGNN, which integrates a tiny graph neural network, combining the speed advantages of the solver with the information benefits of graph learning. It encodes the relationship between past residuals and the parameters $\alpha^{[k]}$ and $\beta$, allowing for more efficient adjustments of the parameters $\alpha$ and $\beta$ to improve the learning capabilities of the IGNN-Solver, as illustrated in Figure 1.
@@
Due to the characteristics of the implicit model—where its representation capability is independent of the forward computation (i.e., the solver has no knowledge of the original task, such as predicting the node’s class, and vice versa)—, we can train this neural solver in a lightweight and unsupervised manner. After the IGNN-Solver is learned, we can utilize it to solve the fixed-point equation of the IGNN layer, thereby further updating the model. The whole training strategy with IGNN-Solver can be found in Appendix B.1 and B.2.
@@
1:frozen IGNN model $f_{\theta}$, initializer $h_{\phi}$, tiny-GNN predictor $s_{\xi}$, sparse subgraph $\hat{{\bm{A}}}_{s}$, storage ${\bm{G}}\in\mathbb{R}^{(m+1)\times d^{\prime}}$.
2:Compute the initial value of fixed point by ${\bm{Z}}^{[0]}=h_{\phi}({\bm{X}})\in\mathbb{R}^{d^{\prime}}$
@@
5:     1) Set $m_{k}=\min\{m,k\}$ and ${\bm{G}}^{[k]}={\bm{G}}\left[0:\left(m_{k}+1\right)\right]\in\mathbb{R}^{\left(m_{k}+1\right)\times d^{\prime}}$
6:     2) Predict ${\alpha}^{[k]},{\beta}^{[k]}=s_{\xi}({\bm{G}}^{[k]},\hat{{\bm{A}}}_{s})$, where ${\alpha}^{[k]}\in\mathbb{R}^{(m_{k}+1)}$, $\mathbf{1}^{\top}{\alpha}^{[k]}=1$ and $\beta^{[k]}\in\mathbb{R}^{1}$
7:     3) Update ${\bm{Z}}^{[k+1]}=\hat{\beta}^{[k]}\cdot\mathbf{1}^{\top}{\bm{G}}^{[k]}+\sum_{i=0}^{m_{k}}\hat{\alpha}_{i}^{[k]}{\bm{Z}}^{[k-m_{k}+i]}$triangleright (Anderson step)
@@
4.1 General Formulation
For a given IGNN layer $f_{\theta}$, input features ${\bm{X}}\in\mathbb{R}^{n\times d}$ and graph $\hat{{\bm{A}}}\in\mathbb{R}^{n\times n}$, we assume that the equation ${\bm{Z}}=f_{\theta}({\bm{Z}},\hat{{\bm{A}}},{\bm{X}})=\sigma(\hat{{\bm{A}}}{\bm{Z}}{\bm{W}}+b_{\Omega}({\bm{X}}))$ has an exact solution at its fixed point. This solution can be obtained using classic solvers (such as Broyden’s method) and running multiple iterations to achieve high precision.
@@ 4.1 General Formulation
The overall structure of the IGNN-Solver is shown in Algorithm 1. Specifically, it learns the parameters $\alpha$ that control the weights among past $m$ steps approximate solution, and $\beta$ that control the weights between past residuals and approximate solution, through a solver $s_{\xi}$ to accelerate the approximate solution of the next update step. To train the IGNN-Solver, we minimize the joint loss function $\mathcal{L}_{\text{total}}$ by back-propagating through this $K$-step temporal process, which is discussed in Section 4.2. It’s worth noting that the original IGNN is frozen (i.e., model parameters $\theta$ are fixed), and only the IGNN-Solver parameters $\xi$ are trained here, so we do not need the ground-truth label $y$ that corresponds to input $x$. This implies that IGNN-Solver can also be fine-tuned during inference after deployment.
4.1.1 Initializer.
@@
To accelerate the convergence of predicted values, we propose constructing an initializer ${\bm{Z}}^{[0]}=h_{\phi}({\bm{X}}):\mathbb{R}^{d}\rightarrow\mathbb{R}^{d^{\prime}}$ for a rapid and reasonable input-based initial estimate value, rather than simply setting the initial values to $\mathbf{0}$ or random, where $\phi$ are parameters of the initializer. We set the intermediate dimension $d^{\prime}$ to be significantly smaller than the feature dimension $d$ to reduce the training overhead of the initializer.
4.1.2 Improved Anderson Iterations with tiny-GNN.
@@
In the original Anderson Acceleration (AA) solver, the parameter $\beta$ is predetermined and fixed [2], whereas the parameters $\alpha^{[k]}$s are determined by solving numerous linear equations using the least squares method. Although this method operates adequately, its efficiency and adaptability in optimizing IGNN models are limited. Therefore, we propose introducing a tiny and learnable graph neural network, as
$\alpha,\beta=s_{\xi}({\bm{G}},\hat{{\bm{A}}})\,\text{, where}\,s_{\xi}:(\mathbb{R}^{(m_{k}+1)\times d^{\prime}}\times\mathbb{R}^{n})\rightarrow(\mathbb{R}^{(m_{k}+1)}\times\mathbb{R}^{1}),$ (4)
@@
to predict the two parameters instead of setting them as the least squares solution on the past residuals ${\bm{G}}$. However, directly using the original graph for modeling would result in an oversized network [53; 59], deviating from the initial goal of using a tiny network for prediction. On the other hand, because of the curse of dimensionality, directly mapping the data from high to extremely low dimensions is not appropriate [27]. Therefore, in response to the challenges posed by large-scale graphs, we took into account both Graph Sparsification and Storage Compression in the design of the solver, as detailed below:
Graph Sparsification.
@@
Firstly, the number of edges in graph $\hat{{\bm{A}}}$ is usually large in practice, as it is affected by the scale of the input (for example, in the ogbn-products dataset, the number of edges in the graph is close to 62M). To maintain the lightweight nature of tiny-GNN and reduce the computational cost of prediction, we propose introducing graph sparsification [24; 36; 55] for a more light graph:
$\hat{{\bm{A}}}_{s}=\varsigma(\hat{{\bm{A}}},\beta),$ (5)
@@
where $\hat{{\bm{A}}}_{s}$ represents the sparse subgraph of $\hat{{\bm{A}}}$, obtained via a graph sparsification operator $\varsigma$, with the parameter $\beta$ controls the sparsity of $\hat{{\bm{A}}}_{s}$. We choose the RPI-Graph approach [53] in this paper, which is a plug-and-play graph sparsification method based on the principle of relevant information. It is worth noting that this subgraph is used only in tiny-GNN predictor $s_{\xi}$, and not in IGNN model $f_{\theta}$.
Storage Compression.
@@
Besides, considering that $n$ (the number of nodes) is relatively large (for example, $n$ is approximately 169K in the ogbn-arxiv dataset), it is not appropriate to map $s_{\xi}$ from a high-dimensional space to a very low-dimensional one in equation 4 directly [45; 57; 44] owing to the curse of dimensionality [27]. Therefore, to maintain $s_{\xi}$ fast and compact, we recommend compress ${\bm{G}}^{[k]}_{i}$ to form a smaller yet still representative version $\bar{{\bm{G}}}^{[k]}_{i}$. Specifically, We map it from $\mathbb{R}^{(m_{k}+1)\times d^{\prime}}$ to an appropriate space $\mathbb{R}^{(m_{k}+1)\times p}$ by multiple layer perception (MLP). Then, the nearest $m_{k}+1$ sets of fixed points are merged into one feature matrix:
@@
where $\mathop{||}_{i=k-m_{k}}^{k}$ represents the concatenation operation that stack the compressed storage matrix $\bar{{\bm{G}}}^{[k]}_{i}\in\mathbb{R}^{(m_{k}+1)\times p}$ in the $k$-th iteration.
Once this smaller but still representative storage matrix $\bar{{\bm{G}}}$ obtained by above, we treat it as a $p$-channel matrix and employ a tiny graph neural network layer
$\alpha^{[k]},\beta^{[k]}=s_{\xi}(\bar{{\bm{G}}}^{[k]},\hat{{\bm{A}}}_{s}),$ (7)
to predict the relative weights $\alpha^{[k]}$ assigned to past $m_{k}+1$ residual, as well as the AA mixing coefficient $\beta^{[k]}$ at the $k$-th iteration.
In this way, $s_{\xi}$ shall autonomously learn and adjust these parameters $\alpha^{[k]}$ and $\beta^{[k]}$ based on the previous solver steps and receiving gradients from subsequent iterations. We provide a comparison of different choices for $s_{\xi}$ in Appendix D, demonstrating that IGNN-Solver improves the convergence path. We also introduce how to train it (see below).
4.2 Training the IGNN-Solver

### v2

@@ III-C Training of IGNNs
Additionally, to train the parameters $\bm{\theta}=\{{\bm{W}},\bm{\Omega}\}$ during the back-propagation process of neural networks, IGNNs compute the Jacobian $\nabla_{\bm{Z}}\mathcal{L}$ by solving the equilibrium equation, as derived from the implicit differential method and the implicit function theorem [42]:
$\nabla_{\bm{Z}}\mathcal{L}={\bm{D}}\odot({\bm{W}}^{\top}\nabla_{\bm{Z}}\mathcal{L}\widehat{{\bm{A}}}^{\top}+\nabla_{\bm{X}}\mathcal{L}),$ (5)
@@
where ${\bm{D}}=\sigma^{\prime}({\bm{W}}{\bm{X}}\widehat{{\bm{A}}}+b_{\Omega}({\bm{U}}))$ and $\sigma^{\prime}(\cdot)$ refers to the element-wise derivative of the map $\sigma$. Once $\nabla_{\bm{Z}}\mathcal{L}$ is obtained, $\nabla_{\bm{W}}\mathcal{L}$ and $\nabla_{\bm{\Omega}}\mathcal{L}$ are derived via the chain rule, eliminating the need to store activations at each layer during gradient updates.
@@
However, IGNNs typically require multiple iterations to reach equilibrium, both in finding the fixed point during the forward process (2) and in solving the implicit differential during the backward process (5), resulting in significant overhead during both training and inference. Therefore, developing a fast and efficient solver to find their fixed points is crucial.
III-D Traditional Anderson Acceleration Solver
Algorithm 1 Anderson Acceleration Solver
3:for $k=0$, $\ldots$, $K-1$ do
6:     3) ${\bm{Z}}^{[k+1]}=\beta\sum_{i=0}^{m_{k}}\alpha_{i}^{[k]}f_{\theta}({\bm{Z}}^{[k-m_{k}+i]})+(1-\beta)\sum_{i=0}^{m_{k}}\alpha_{i}^{[k]}{\bm{Z}}^{[k-m_{k}+i]}\quad$triangleright Anderson step
7:end for
Since implicit networks typically require multiple iterations to reach equilibrium, a commonly used solver method is Anderson Acceleration [13], which is widely applied in [3, 15, 14, 16]. Specifically, Given an implicit network layer $f_{\theta}$ and input ${\bm{X}}$, we aim to find the fixed-point ${\bm{Z}}^{*}$ in the system that represents the equilibrium state of the network output by solving a root-finding problem
$g_{\theta}\left({\bm{Z}}^{\star},{\bm{X}}\right):=f_{\theta}\left({\bm{Z}}^{\star},{\bm{X}}\right)-{\bm{Z}}^{\star}=0,$ (6)
via utilizing fixed-point solvers, such as the classical Anderson Acceleration (AA) [13], which can directly seek the equilibrium point ${\bm{Z}}^{\star}$ through quasi-Newton methods. This approach demonstrates super-linear convergence properties.
We briefly introduce the AA-solver, as our approach is inspired by it. Algorithm 1 provides a pseudocode example, where ${\bm{G}}^{[k]}=[g_{\theta}(\mathbf{z}^{[k-m_{k}]}),\ldots,g_{\theta}(\mathbf{z}^{[k]})]$ represents the past residuals. The key idea behind AA is to accelerate the next approximate solution by leveraging information stored from the history $m$ iteration steps, constructing a normalized linear combination of past updates with weights $\alpha^{[k]}$ and $\beta$. These weights are computed greedily to minimize the linear combination of past AA update steps. This approach is typically effective for accelerating the solution of implicit functions.
However, since the computation process relies only on the final output and does not require storing any intermediate activation values [14], the memory consumption during the training process remains constant, which is also a key reason why fixed-point solvers are attractive.
Figure 1: Comparison of our IGNN-Solver and the generic Anderson solver at each iteration. (a) The canonical Anderson solver performs local least-squares fitting at each iteration, with a fixed $\beta=\beta^{[k]}$ set to a constant. (b) Our learnable IGNN-Solver improves and incorporates learnable iterative updates for enhanced convergence via tiny GNN.
IV Our purposed IGNN-Solver
@@
Although traditional fixed-point solvers for IGNNs demonstrate functionality, they are characterized by relatively slow performance, necessitate manual parameter tuning, and are not well-suited for graph-based applications. To address these limitations, we propose a lightweight, learnable, and content-aware fixed-point solver for IGNN that incorporates a tiny graph neural network, combining the speed advantages of the solver with the information benefits of graph learning. By encoding the relationship between past residuals and parameters $\alpha^{[k]}$ and $\beta$, it enables more efficient adjustments to enhance the learning capabilities of the IGNN-Solver, as shown in Figure 1.
@@
Given the implicit model’s characteristic that its representation capability is decoupled from the forward computation (i.e., the solver has no knowledge of the task, such as node classification, and vice versa), we can train this neural solver in a lightweight, unsupervised manner. Once learned, the IGNN-Solver can solve the fixed-point equation within the IGNN layer, facilitating further model updates. Note that our solver does not alter the original structure of the model but only accelerates the solution process. Details of the training strategy involving the IGNN-Solver are provided in Appendix B.
@@
1:frozen IGNN model $f_{\theta}$, initializer $h_{\phi}$, tiny-GNN predictor $s_{\xi}$, sparse subgraph $\widehat{{\bm{A}}}_{s}$, storage ${\bm{G}}\in\mathbb{R}^{(m+1)\times d^{\prime}}$.
2:Compute the initial value of fixed point by ${\bm{Z}}^{[0]}=h_{\phi}({\bm{X}})\in\mathbb{R}^{d^{\prime}}$triangleright initializer to guess initial value
@@
5:     1) Set $m_{k}=\min\{m,k\}$ and ${\bm{G}}^{[k]}={\bm{G}}\left[0:\left(m_{k}+1\right)\right]\in\mathbb{R}^{\left(m_{k}+1\right)\times d^{\prime}}$triangleright compress storage
6:     2) Predict ${\alpha}^{[k]},{\beta}^{[k]}=s_{\xi}({\bm{G}}^{[k]},\widehat{{\bm{A}}}_{s})$, where ${\alpha}^{[k]}\in\mathbb{R}^{(m_{k}+1)}$, $\mathbf{1}^{\top}{\alpha}^{[k]}=1$ and $\beta^{[k]}\in\mathbb{R}^{1}$triangleright predict with tiny-GNN
7:     3) Update ${\bm{Z}}^{[k+1]}=\beta^{[k]}\cdot\mathbf{1}^{\top}{\bm{G}}^{[k]}+(1-\beta^{[k]})\sum_{i=0}^{m_{k}}\alpha_{i}^{[k]}{\bm{Z}}^{[k-m_{k}+i]}$triangleright Anderson step
@@
IV-A General Formulation
Let ${\bm{X}}\in\mathbb{R}^{d\times n}$ be the input features and $\widehat{{\bm{A}}}\in\mathbb{R}^{n\times n}$ be the graph, IGNN learns the node representation by finding the fixed point below
${\bm{Z}}=f_{\theta}({\bm{Z}},\widehat{{\bm{A}}},{\bm{X}})=\sigma({\bm{W}}{\bm{Z}}\widehat{{\bm{A}}}+b_{\Omega}({\bm{X}}))$ (7)
@@ IV-A General Formulation
The overall structure of the IGNN-Solver is shown in Algorithm 2. Specifically, it learns the parameters $\alpha$ that control the weights among past $m$ steps approximate solution, and $\beta$ that control the weights between past residuals and approximate solution, through a solver $s_{\xi}$ to accelerate the approximate solution of the next update step. To train the IGNN-Solver, we minimize the joint loss function $\mathcal{L}_{\text{total}}$ by back-propagating through this $K$-step temporal process, which is discussed in Section IV-C. It’s worth noting that the original IGNN is frozen (i.e., model parameters $\theta$ are fixed), and only the IGNN-Solver parameters $\xi$ are trained here, so we do not need the ground-truth label $y$ that corresponds to input $x$. This implies that IGNN-Solver can also be fine-tuned during inference after deployment.
IV-A1 Initializer
To accelerate the convergence of predicted values, we propose constructing an initializer
${\bm{Z}}^{[0]}=h_{\phi}({\bm{X}})$ (8)
@@ IV-A1 Initializer
where $h_{\phi}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d^{\prime}}$, for a rapid and reasonable input-based initial estimate value, rather than simply setting the initial values to $\mathbf{0}$ or random, where $\phi$ are learnable parameters of the initializer. We set the intermediate dimension $d^{\prime}$ to be significantly smaller than the feature dimension $d$ to reduce the training overhead of the initializer.
IV-A2 Improved Anderson Iterations with tiny-GNN
@@
In the original AA solver, the parameter $\beta$ is predetermined and fixed [13], whereas the parameters $\alpha^{[k]}$s are determined by solving numerous linear equations using the least squares method. Although this method operates adequately, its efficiency and adaptability in optimizing IGNN models are limited. Therefore, we propose introducing a tiny and learnable graph neural network as
$\alpha,\beta=s_{\xi}({\bm{G}},\widehat{{\bm{A}}})\,$ (9)
where $\,s_{\xi}:(\mathbb{R}^{(m_{k}+1)\times d^{\prime}}\times\mathbb{R}^{n})\rightarrow(\mathbb{R}^{(m_{k}+1)}\times\mathbb{R}^{1})$, to predict the two parameters instead of setting them as the least squares solution on the past residuals ${\bm{G}}$.
Remark 1 (Well-posedness of IGNN-Solver).
The proposed IGNN-Solver remains well-posed when the IGNN satisfies the reparameterization defined in (4). Note that the IGNN-Solver introduces two key improvements while maintaining the structure and parameters of the model. First, we propose an initializer for a better initial guess, establishing a more meaningful starting point, reducing solving time, and shortening the fixed-point trajectory. Second, we parameterize the Anderson weights and step sizes within their original range, ensuring that well-posedness is preserved. Consequently, the modifications introduced in the IGNN-Solver retain the foundational characteristics of classical AA without compromising the well-posedness condition. Therefore, for an IGNN $(f_{\theta},b_{\Omega},\sigma)$ where $f_{\theta}$ satisfies the well-posedness condition, it holds that the corresponding IGNN-Solver $(h_{\phi},s_{\xi})$ on this IGNN is also well-posed, thereby guaranteeing the existence and uniqueness of its solution.
IV-B Graph Sparsification and Storage Compression
@@
Directly using the original graph for modeling would result in an oversized network [17, 18], deviating from the initial goal of using a tiny network for prediction. Moreover, because of the curse of dimensionality, directly mapping the data from high to extremely low dimensions is not appropriate [19]. Therefore, in response to the challenges posed by large-scale graphs, we take into account both Graph Sparsification and Storage Compression in the design of the solver, as detailed below.
IV-B1 Graph Sparsification
@@
Firstly, the number of edges in graph $\widehat{{\bm{A}}}$ is usually large in practice, as it is affected by the scale of the input (for example, in the ogbn-products dataset, the number of edges in the graph is close to 62M). To maintain the lightweight nature of tiny-GNN and reduce the computational cost of prediction, we propose introducing graph sparsification [43, 44, 45] for a light graph:
$\widehat{{\bm{A}}}_{s}=\varsigma(\widehat{{\bm{A}}},\beta),$ (10)
@@
where $\widehat{{\bm{A}}}_{s}$ represents the sparse subgraph of $\widehat{{\bm{A}}}$, obtained via a graph sparsification operator $\varsigma$, with the parameter $\beta$ controls the sparsity of $\widehat{{\bm{A}}}_{s}$. We choose the RPI-Graph approach [17] in this paper, which is a plug-and-play graph sparsification method based on the principle of relevant information. It is worth noting that this subgraph is used exclusively in the tiny-GNN predictor $s_{\xi}$ and does not alter the original graph of the IGNN model $f_{\theta}$.
IV-B2 Storage Compression
@@
Besides, considering that $n$ (the number of nodes) is relatively large (for example, $n$ is approximately 169K in the ogbn-arxiv dataset), it is not appropriate to map $s_{\xi}$ from a high-dimensional space to a very low-dimensional one in (9) directly [25, 46, 47] owing to the curse of dimensionality [19]. Therefore, to maintain $s_{\xi}$ fast and compact, we recommend compress ${\bm{G}}^{[k]}_{i}$ to form a smaller yet still representative version $\bar{{\bm{G}}}^{[k]}_{i}$. Specifically, We map it from $\mathbb{R}^{(m_{k}+1)\times d^{\prime}}$ to an appropriate space $\mathbb{R}^{(m_{k}+1)\times p}$ by multiple layer perception (MLP). Then, the nearest $m_{k}+1$ sets of fixed points are merged into one feature matrix:
@@
where $\mathop{||}_{i=k-m_{k}}^{k}$ represents the concatenation operation that stacks the compressed storage matrix $\bar{{\bm{G}}}^{[k]}_{i}\in\mathbb{R}^{(m_{k}+1)\times p}$ in the $k$-th iteration.
Once the smaller yet representative storage matrix $\bar{{\bm{G}}}$ is obtained as described above, we treat it as a $p$-channel matrix and employ a tiny graph neural network layer:
$\alpha^{[k]},\beta^{[k]}=s_{\xi}(\bar{{\bm{G}}}^{[k]},\widehat{{\bm{A}}}_{s}),$ (12)
to predict the relative weights $\alpha^{[k]}$ assigned to past $m_{k}+1$ residual, as well as the AA mixing coefficient $\beta^{[k]}$ at the $k$-th iteration. Notably, we normalize the original output $\widehat{\alpha}^{[k]}$ using the function
$\alpha^{[k]}=\widehat{\alpha}^{[k]}+\frac{\left(1-\mathbf{1}^{\top}\widehat{\alpha}^{[k]}\right)}{m_{k}+1}\cdot\mathbf{1}$
to ensure that the constraint $\mathbf{1}^{\top}\alpha^{[k]}=1$ is satisfied.
In this way, $s_{\xi}$ shall autonomously learn and adjust these parameters $\alpha^{[k]}$ and $\beta^{[k]}$ based on previous solver steps, while also receiving gradients from subsequent iterations. We provide a comparison of different choices for $s_{\xi}$ in Section V-D, demonstrating that IGNN-Solver improves the convergence path. We also introduce the training procedure below.
IV-C Training IGNN-Solver